40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Dec 14 14:10:36 2021, aaron, Summary, PSOMA, PSOMA calibration parameters, minor modifications 
    Reply  Fri Dec 17 12:46:41 2021, aaron, Summary, PSOMA, PSOMA calibration parameters, minor modifications 
       Reply  Fri Dec 17 18:46:55 2021, aaron, Summary, PSOMA, PSOMA calibration parameters, minor modifications 
Message ID: 2836     Entry time: Fri Dec 17 12:46:41 2021     In reply to: 2834     Reply to this: 2837
Author: aaron 
Type: Summary 
Category: PSOMA 
Subject: PSOMA calibration parameters, minor modifications 

I removed the 10 dB attenuator between the LB box output and the laser driver's current modulation input. The ranges of the LB box output and ITC 502 input are matched, so there should be no need for an attenuator. I also slightly reduced the servo gain to 5.3. There is still a 10dB attenuator between the PDH mixer/lowpass and the LB box input, to avoid saturating the servo's input stage (+- 330 mV).

After moving most electronics to the same power strip as in the above elog and tying the table to ground, I no longer see a 144 kHz oscillation in the PDH error signal (oscilloscope). I do see a ~40 Hz oscillation in the PDH error signal... could that be the fundamental mode of the cantilever? Zach's thesis quotes a fundamental resonance at 45 Hz. The 40 Hz oscillation is present even while unlocked (but with the servo engaged).

Though the error signal looked clean, the lock was not steady for more than 10s of seconds. I suspected that the current control loop was far from the center of its range. I am monitoring the current control signal from the LB box Aux output, which picks off the signal before the final summing amplifier stage. Therefore, in order to adjust the temperature setpoint to 'zero' the control signal, I needed to first disconnect the LB box inputs then adjust the output offset until the voltage at aux output matched that at the final output. Then, with the laser locked, I adjusted the temperature setpoint until the control signal as measured at aux output was zero (plus some offset that was present at both the main output and aux output even with the inputs disconnected).

After adjusting the temperature setpoint manually, the beamspot is noticeably brighter on the trans mon camera, and the lock is stable for minutes.

I tried setting up the digital control to handle the temperature loop, but the realtime system is in a bad state and I didn't attempt to diagnose it.

ELOG V3.1.3-